How would you prepare a 50mL buffer at pH = 8.90 given the following materials: 100mL volumetric flask, 100mL beaker, an unlimited supply of deionized water, 1.0M NH3, 0.50M NH4Cl, 0.15M CH3COOH, 0.10M NaOH, 0.10M HCl, a stir bar, a calibrated pH probe, a burette and a lab quest. include relevant volumes/masses, type of glassware used, significant figures and units.
Answer #1

To form buffer of pH 8.90, ammonia and NH4Cl is to be used.

pOH of required buffer = 14 - 8.90 = 5.10

Kb of ammonia = 1.8 x 10-5

pKb = -log(1.8 x 10-5) = 4.75

According Henderson's equation for weak base and its salt

pOH = pKb + log[NH4Cl]/[NH3]

5.10 = 4.75 + log[NH4Cl]/[NH3]

100.35 = 2.24 = [NH4Cl]/[NH3]. (Equation 1)

Hence, ratio of [NH4Cl] : [NH3] is 2.24 : 1

To get that ratio of concentrations, let the volume of 0.50 M NH4Cl to be added is V ml

Required volume of buffer = 50 ml

Hence, effective concentration of NH4Cl is (V x 0.50)/50

Volume of 1.0 M NH3 = (50 -V ml)

Effective concentration of NH3 = {(50-V) x 1.0}/50

Substituting the values of effective concentrations in equation 1

2.24 = (V x 0.50)/{(50-V) x 1.0}

2.24(50-V) = 0.50 V

112 - 2.24 V = 0.50 V

V = 112/2.75 = 40.9 ml

So, 40.9 ml of 0.50 M NH4Cl solution should be taken in a 100 mL volumetric flask using a burette.

To it (50-40.9) i.e. 9.1 ml of 1.0 M NH3 solution should be added using another burette.

Stir the solutions and stopper the flask and label the solution as buffer of pH = 8.90